A standing wave is set up in a string 120 cm long. The string vibrates in four segments under 

a frequency of 120 Hz. Calculate the wavelength and the fundamental frequency of the string


1
Expert's answer
2021-05-16T17:53:30-0400

"L =120\ cm =1.2 \ m"

"n = 4 \"

"\lambda= \frac{2}{n}*L= \frac{2}{4}*1.2=0.6"

"f = 120"

"f= f_n*n;f_n= \frac{f}{n}= \frac{120}{4}=3"

"\text {Answer : }\lambda = 0.6\ m ; f_4= 30\ Hz"